Hello, I have thought about this one for a couple of days and don't even know where to begin. Any assistance would be greatly appreciated!


If the product of the integers w, x, y, and z is 770, and if 1 < w < x < y < z, what is w + z.
a) 10
b) 13
c) 16
d) 18
e) 21

Answer by askmemath(368)

Split 770 into its factors
770 = 77x10
77x10 = 7x11 x 5x2
so w,x,y,z = 2,5,7 ,11 respectively
and w+z = 2+11 = 13